OpenStudy (anonymous):

3p-2/2=p-6,then p=??

OpenStudy (lgbasallote):

cross multiply first 3p - 2 = 2p - 12 isolate p 3p - 2p = -12 + 10 p = -2

OpenStudy (anonymous):

thanks igbasallote

OpenStudy (mayliz):

p=-5/2

OpenStudy (anonymous):

ok :)